# FineMolTex: Towards Fine-grained Molecular Graph-Text Pre-training

### v2

@@ 5 Conclusions, Limitations and Broader Impact
In this paper, we reveal that fine-grained motif-level knowledge is crucial for molecular representation learning. We propose FineMolTex to jointly learn both coarse- and fine-grained knowledge through a contrastive alignment task and a masked multimodal learning task, respectively. By masking the fine-grained tokens and predicting their labels using tokens from the other modality, we can effectively learn fine-grained alignment between motifs and words. Experimental results on three downstream tasks and two case studies demonstrate the effectiveness of FineMolTex.
Limitations. Despite the promising results of FineMolTex, the molecular graph-text dataset is small in scale, hindering the cross-modal understanding. To overcome the limitation, we plan to seek more high-quality texts closely related to molecular graphs or devise an approach to align the embedding spaces of molecular graphs and texts without paired data.
Broader Impact. We demonstrate the importance of capturing fine-grained motif-level knowledge through diverse downstream tasks. Consequently, our research has the potential to contribute to the development of a foundation model for various chemical and biomedical applications, paving a promising way for key scientific inquiries, including drug discovery, molecular reaction prediction, and molecular property prediction.
@@ Appendix A Related Work
Molecular Learning Based on Graphs. Molecular graphs are commonly employed as inputs for molecular learning. Numerous studies employ various GNNs as base encoders to facilitate molecular representation learning for downstream tasks [48, 42]. These approaches generally require supervised signals and are not applicable to other tasks. Recent advancements in research have introduced self-supervised learning methods. Some methods focus on reconstruction tasks for pre-training. PreGNN [13] enhances GNN pre-training through context prediction and node/edge attribute masking. GROVER [33] introduces molecular-specific self-supervised techniques, including contextual property prediction and graph-level motif prediction. MGSSL [55] adopts a motif-based graph self-supervised strategy that predicts the topology and labels of motifs. Additionally, some methods leverage contrastive learning for pertaining, and general graph augmentation methods [50] are also applicable to molecular datasets. InfoGraph [36] optimizes model training by maximizing the mutual information between the global graph representations and their substructures of varying granularity. Incorporating chemical domain knowledge, MoCL [37] employs two novel molecular graph augmentation methods: replacing specific substructures or altering a few carbon atoms. MolR [40] aims to maintain the equivalence relation between reactants and products within the embedding space. Despite the effectiveness of pre-trained models, it is still challenging to generalize new categories or tasks without labeled examples or fine-tuning.
@@
Fine-grained Vision-Language models. Various works design fine-grained alignment methods to better align the region visual features (RoIs) into text features of the vision-language model (VLM). OV-DETR [52] introduces a transformer-based approach for open vocabulary object detection, using conditional binary matching instead of traditional bipartite matching. VLDet [23] aligns objects and language by transforming images into regions and captions into words, employing a set matching strategy. DetCLIPv2 [49] leverages ATSS [54] for object detection, trained on a standard detection dataset, a grounding dataset, and an image-text pairs dataset. BARON [45] aligns embeddings within groups of related regions, processing them through a text encoder. CoDet [28] treats region-word alignment as a co-occurring object discovery problem. F-VLM [19] uses a CLIP vision encoder and a VLM feature pooler for region features, combining detection scores with VLM predictions. RO-ViT [15] addresses positional embedding gaps in vision-language pre-training by introducing a cropped positional embedding module, enhancing alignment for downstream tasks. While these methods rely on a detection model to assign labels to the RoI, our approach lacks such a model to provide supervised signals for motifs, making it challenging to align fine-grained information.
@@ Appendix B Technical Details of Text-based Molecule Editing Task
Following the molecule editing framework proposed by [26], we utilize FineMolTex and the generation module, including an encoder and a decoder, to generate molecules. This process is structured into two phases. The first phase is the space alignment, which utilizes two projectors, $p_{m}$ and $p_{g}$, to align the representation space of the generative model with our model into a joint representation space following a contrastive learning strategy as:
$L_{\text{ali}}=$ $-\dfrac{1}{2}\mathbb{E}_{m}[\log\dfrac{\exp(\cos(\mathbf{z_{m_{0}}},p_{g}(\mathbf{w}))/\tau)}{\exp(\cos(\mathbf{z_{m_{0}}},p_{g}(\mathbf{w}))/\tau)+\sum_{w^{\prime}}\exp(\cos(\mathbf{z_{m_{0}}},p_{g}(\mathbf{w^{\prime}}))/\tau)}$ (11)
$+\log\dfrac{\exp(\cos(w,p_{m}(\mathbf{z_{m_{0}}}))/\tau)}{\exp(\cos(\mathbf{w},p_{m}(\mathbf{z_{m_{0}}}))/\tau)+\sum_{\mathbf{z_{m_{0}^{\prime}}}}\exp(\cos(\mathbf{w},p_{m}(\mathbf{z_{m_{0}^{\prime}}}))/\tau)}],$
@@
$\mathbf{w^{*}}=\text{argmin}_{\mathbf{w^{*}}}(cos(p_{g}(\mathbf{w^{*}}),\mathbf{z_{t_{0}}})/\tau)+\lambda l_{2}(\mathbf{w^{*}},\mathbf{w}).$ (12)
@@ Appendix C Reproducibility Information
C.1 Dataset Statistics
C.1.1 Pre-training Dataset
@@
We employ a dataset known as PubChemSTM [26], derived from the PubChem database [16], which consists of chemical structure-text pairs. The dataset is available in two versions: PubChemSTM-raw, which retains the original annotations, and PubChemSTM-extracted, where the names of the molecules are replaced with the generic term "this molecule". In our study, we use the PubChemSTM-extracted version. Examples of PubChemSTM-extracted are illustrated in Figure  6.
Figure 6: Examples on PubChemSTM-extracted
C.1.2 Retrieval Datasets
@@
We utilize three pertinent fields from the DrugBank database [44] for each small molecule drug in our zero-shot retrieval task: the Description field, the Pharmacodynamics field, and the Anatomical Therapeutic Chemical (ATC) classification. Specifically, the DrugBank-Description provides a comprehensive overview of the drug’s chemical characteristics, historical development, and regulatory status. The DrugBank-Pharmacodynamics section details the mechanisms through which the drug influences or alters the organism in which it is used. This includes both desired and undesired effects (commonly referred to as side effects). Lastly, the DrugBank-ATC classification system organizes the molecules into groups based on the organs or systems they target, along with their therapeutic, pharmacological, and chemical properties. The datasets consist of 1154, 1005, and 3007 molecule-text pairs for each field, respectively.
C.1.3 Molecule Property Prediction Datasets
@@
The MoleculeNet dataset, used for molecular property prediction in downstream tasks, contains two main categories. Some datasets are used for pharmacological property prediction. The Blood-Brain Barrier Penetration (BBBP) [30] dataset evaluates whether a molecule can penetrate the central nervous system. The three datasets related to toxicity—Tox21 [5], ToxCast [46], and ClinTox [10]—assess the toxicity of molecular compounds. The Side Effect Resource (SIDER) [18] dataset contains information on adverse drug reactions from a marketed drug database. Other datasets are used for biophysical property prediction. The Maximum Unbiased Validation (MUV) [32] dataset, a subset of the PubChem BioAssay (PCBA), is created using a refined nearest neighbor analysis. The HIV dataset, sourced from the Drug Therapeutics Program (DTP) AIDS Antiviral Screen [51], focuses on predicting the inhibition of HIV replication. The BACE dataset, included in MoleculeNet [46], measures the binding affinity of various inhibitors to $\beta$-secretase 1 (BACE-1). The overall dataset statistics of MoleculeNet are shown in Table  5.
@@
Table 5: Dataset statistics of MoleculeNet for molecule property prediction task.
@@
C.2 Baselines
@@ C.3 Operating Environment
C.4 Inplementation Details
We use Pytorch to implement our model. For the pre-trained baselines, we directly utilize the checkpoints provided by the authors. For other baselines, we utilize the original codes from their authors and train the models in an end-to-end way. We utilize the BRICS [6] to fragment molecules into motifs. As BRICS primarily cleaves bonds according to a predefined set of chemical reactions, often resulting in several large fragments per molecule, we further utilize the post-processing procedure proposed by MGSSL [55], which is designed to minimize the number of ring variants and facilitate the cleavage of side chains. Subsequently, we build a vocabulary of all motif tokens identified in the PubChemSTM dataset, which comprises a total of 30,080 unique motifs. We note that within our dataset, certain motifs are infrequently present, while others are prevalent but lack significant semantic value. This uneven distribution presents a challenge for the model, as it struggles to extract meaningful insights from such disparate occurrences. Thus we have constructed a masking set of 2,457 motifs that excludes motifs appearing fewer than 8 times or more than 80,005 times. Based on this masking set, we selectively mask motifs to enhance the model’s learning efficiency and focus on extracting valuable information from the most informative tokens. To map the embedding of the molecule and the text into the same dimension, we further utilize a projector MLP layer for the graph encoder. We utilize Adam as the optimizer for the GNN encoder, the language model, the two projector MLPs, the multi-modal framework, and the motif classifier with different learning rates, and test the learning rate ranging from {1e-3, 1e-4, 1e-5}. We set $\ell=2$, and test $\ell_{\text{trm}_{M}}$ raging from {1,2,3,4,5}, $\ell_{\text{trm}_{T}}$ ranging from {8,10,12}. For the coefficient of loss function, we test $\alpha$ and $\beta$ raging from {0.5,1,2}. We train our model with a total batch size of 16 for 15 epochs. For fair comparisons, we randomly run 5 times and report the average results for all methods.
Optimal Parameters. The optimal parameters can be found in the config.json file in our code repository. Specifically, we set the dimension of the graph encoder to 300 and the dimensions of the transformer and cross-attention layer to 768. The hyperparameters are set as $\alpha=0.5$ and $\beta=1$. For $\ell=2$, we set $\ell_{\text{trm}M}$ to 2 for the first round and 3 for the second round, and $\ell{\text{trm}_{T}}$ to 10 for the first round and 12 for the second round. The learning rates are set as follows: 1e-5 for the graph encoder, text encoder, transformer layers, and cross-attention layer; 3e-5 for the projector MLP of the graph encoder; and 1e-3 for the motif classifier.
@@ Appendix D Additional Resuts
D.1 Graph-text Retrieval Task on Drugbank-Description
The results of the graph-text retrieval task on Drugbank-Description can be found in Table  LABEL:tab:des. We can observe that our model consistently performs better than baselines.
Table 6: Accuracy (%±$\pm\sigma$) of graph-text retrieval task on DrugBank-Description.
Given Molecular Graph Given Text
T 4 10 20 4 10 20
FineMolTex 99.58±$\mathbf{\pm 0.05}$ 97.97±$\mathbf{\pm 0.00}$ 96.45±$\mathbf{\pm 0.16}$ 99.58±$\mathbf{\pm 0.04}$ 97.89±$\mathbf{\pm 0.08}$ 96.11±$\mathbf{\pm 0.12}$
Figure 7: Visualization of motif tokens and word tokens using t-SNE with legend.
D.2 Visual Analysis of the Output Molecules on 4 Text-based Molecule Editing Tasks
We present additional visual analyses of the output molecules for four text-based molecule editing tasks, as depicted in Figure  8. FineMolTex effectively generates motifs that align with expectations, demonstrating its superior ability to learn fine-grained knowledge.
Figure 8: Additional visual analysis of the output molecules of MoleculeSTM and Motif-MolTex on 4 text-based molecule editing tasks. Differences between the input molecule and output molecule of FineMolTex are highlighted in red and green circles. Higher LogP indicates lower water solubility.
D.3 Visualization of Motif and Word Tokens with Legends
Figure  7 displays the complete visualization with legends, illustrating that the embeddings of relevant motifs and words are closer in the embedding space.
D.4 Explanation of Predictions on Masked Motifs Based on Word Tokens
We provide additional explanations for the predictions of FineMolTex in Figure  9, verifying whether the words with the highest interpretive weights are relevant to the masked motifs. In (1), "indanone" is directly the name of the masked motif. In (2), the masked motif is part of "azobenzene". In (3), "azetidine" is part of the masked motif. In (4), the masked motif is a substructure of "trimethylamino" and "butyrobetaine". These explanations demonstrate that FineMolTex has learned fine-grained knowledge, providing insights for the prediction task.
Figure 9: Explanations of prediction of motifs based on word tokens.

### v5

@@ 5. Conclusions
In this paper, we reveal that fine-grained motif-level knowledge is crucial for molecular representation learning. We propose FineMolTex to jointly learn both coarse- and fine-grained knowledge through a contrastive alignment task and a masked multimodal learning task, respectively. By selectively masking the important motif/word tokens and predicting their labels using tokens from the other modality, we can effectively learn fine-grained alignment between motifs and words. Experimental results on three downstream tasks and two case studies demonstrate the effectiveness of FineMolTex.
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China (No. 62192784, U22B2038, 62472329). This research is also supported by the Ministry of Education, Singapore under its Academic Research Fund (AcRF) Tier 1 grant (22-SIS-SMU-054). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
@@ Appendix A Related Work
Molecular Learning Based on Graphs. Molecular graphs are commonly employed as inputs for molecular learning. Numerous studies employ various GNNs as base encoders to facilitate molecular representation learning for downstream tasks (Yang et al., 2021; Wang et al., 2021a). These approaches generally require supervised signals and are not applicable to other tasks. Recent advancements in research have introduced self-supervised learning methods. Some methods focus on reconstruction tasks for pre-training. PreGNN (Hu et al., 2020) enhances GNN pre-training through context prediction and node/edge attribute masking. GROVER (Rong et al., 2020) introduces molecular-specific self-supervised techniques, including contextual property prediction and graph-level motif prediction. MGSSL (Zhang et al., 2021) adopts a motif-based graph self-supervised strategy that predicts the topology and labels of motifs. Additionally, some methods leverage contrastive learning for pertaining, and general graph augmentation methods (You et al., 2020) are also applicable to molecular datasets. InfoGraph (Sun et al., 2019) optimizes model training by maximizing the mutual information between the global graph representations and their substructures of varying granularity. Incorporating chemical domain knowledge, MoCL (Sun et al., 2021) employs two novel molecular graph augmentation methods: replacing specific substructures or altering a few carbon atoms. MolR (Wang et al., 2022) aims to maintain the equivalence relation between reactants and products within the embedding space. Despite the effectiveness of pre-trained models, it is still challenging to generalize new categories or tasks without labeled examples or fine-tuning.
@@
Fine-grained Vision-Language models. Various works design fine-grained alignment methods to better align the region visual features (RoIs) into text features of the vision-language model (VLM). OV-DETR (Zang et al., 2022) introduces a transformer-based approach for open vocabulary object detection, using conditional binary matching instead of traditional bipartite matching. VLDet (Lin et al., 2022) aligns objects and language by transforming images into regions and captions into words, employing a set matching strategy. DetCLIPv2 (Yao et al., [n. d.]) leverages ATSS (Zhang et al., 2020) for object detection, trained on a standard detection dataset, a grounding dataset, and an image-text pairs dataset. BARON (Wu et al., [n. d.]) aligns embeddings within groups of related regions, processing them through a text encoder. CoDet (Ma et al., 2024) treats region-word alignment as a co-occurring object discovery problem. F-VLM (Kuo et al., 2022) uses a CLIP vision encoder and a VLM feature pooler for region features, combining detection scores with VLM predictions. RO-ViT (Kim et al., [n. d.]) addresses positional embedding gaps in vision-language pre-training by introducing a cropped positional embedding module, enhancing alignment for downstream tasks. While these methods rely on a detection model to assign labels to the RoI, our approach lacks such a model to provide supervised signals for motifs, making it challenging to align fine-grained information.
Graph Neural Networks. Graph Neural Networks (GNNs) are powerful deep learning algorithms that can be used to model graph-structured data. In recent years, there have been enormous successful applications of GNNs on various areas such as social media mining (Yu et al., 2024a, b, 2025; Lei et al., 2022), graph cluster (Liu et al., 2022, 2023b, 2024), drug discovery (Jin et al., 2018), and recommender system (Berg et al., 2017; Ying et al., 2018).
@@ Appendix B Technical Details of Text-based Molecule Editing Task
Following the molecule editing framework proposed by (Liu et al., 2023c), we utilize FineMolTex and the generation module, including an encoder and a decoder, to generate molecules. This process is structured into two phases. The first phase is the space alignment, which utilizes two projectors, $p_{m}$ and $p_{g}$, to align the representation space of the generative model with our model into a joint representation space following a contrastive learning strategy as:
(13) $L_{\text{ali}}=$ $-\dfrac{1}{2}\mathbb{E}_{m}[\log\textstyle\frac{\exp(\cos(\mathbf{z_{m_{0}}},p_{g}(\mathbf{w}))/\tau)}{\exp(\cos(\mathbf{z_{m_{0}}},p_{g}(\mathbf{w}))/\tau)+\sum_{w^{\prime}}\exp(\cos(\mathbf{z_{m_{0}}},p_{g}(\mathbf{w^{\prime}}))/\tau)}$
$+\textstyle\log\frac{\exp(\cos(w,p_{m}(\mathbf{z_{m_{0}}}))/\tau)}{\exp(\cos(\mathbf{w},p_{m}(\mathbf{z_{m_{0}}}))/\tau)+\sum_{\mathbf{z_{m_{0}^{\prime}}}}\exp(\cos(\mathbf{w},p_{m}(\mathbf{z_{m_{0}^{\prime}}}))/\tau)}],$
@@
(14) $\mathbf{w^{*}}=\text{argmin}_{\mathbf{w^{*}}}(cos(p_{g}(\mathbf{w^{*}}),\mathbf{z_{t_{0}}})/\tau)+\lambda l_{2}(\mathbf{w^{*}},\mathbf{w}).$
@@ Appendix C Reproducibility Information
C.1. Dataset Statistics
C.1.1. Pre-training Dataset
@@
We employ a dataset known as PubChemSTM (Liu et al., 2023c), derived from the PubChem database (Kim et al., 2021), which consists of chemical structure-text pairs. The dataset is available in two versions: PubChemSTM-raw, which retains the original annotations, and PubChemSTM-extracted, where the names of the molecules are replaced with the generic term ”this molecule”. In our study, we use the PubChemSTM-extracted version. Examples of PubChemSTM-extracted are illustrated in Figure  7.
Figure 7. Examples on PubChemSTM-extracted
C.1.2. Retrieval Datasets
@@
We utilize three pertinent fields from the DrugBank database (Wishart et al., 2018) for each small molecule drug in our zero-shot retrieval task: the Description field, the Pharmacodynamics field, and the Anatomical Therapeutic Chemical (ATC) classification. Specifically, the DrugBank-Description provides a comprehensive overview of the drug’s chemical characteristics, historical development, and regulatory status. The DrugBank-Pharmacodynamics section details the mechanisms through which the drug influences or alters the organism in which it is used. This includes both desired and undesired effects (commonly referred to as side effects). Lastly, the DrugBank-ATC classification system organizes the molecules into groups based on the organs or systems they target, along with their therapeutic, pharmacological, and chemical properties. The datasets consist of 1154, 1005, and 3007 molecule-text pairs for each field, respectively. To test the generalizability of FineMolTex on unseen molecules and texts, given a molecular graph, we select its corresponding molecular description as the positive candidate, and randomly select molecular descriptions from other molecules as negative candidates, aiming to identify the textual description that best aligns with the molecular graph, following (Liu et al., 2023c).
C.1.3. Molecule Property Prediction Datasets
@@
The MoleculeNet dataset, used for molecular property prediction in downstream tasks, contains two main categories. Some datasets are used for pharmacological property prediction. The Blood-Brain Barrier Penetration (BBBP) (Martins et al., 2012) dataset evaluates whether a molecule can penetrate the central nervous system. The three datasets related to toxicity—Tox21 (Challenge, 2014), ToxCast (Wu et al., 2018), and ClinTox (Gayvert et al., 2016)—assess the toxicity of molecular compounds. The Side Effect Resource (SIDER) (Kuhn et al., 2016) dataset contains information on adverse drug reactions from a marketed drug database. Other datasets are used for biophysical property prediction. The Maximum Unbiased Validation (MUV) (Rohrer and Baumann, 2009) dataset, a subset of the PubChem BioAssay (PCBA), is created using a refined nearest neighbor analysis. The HIV dataset, sourced from the Drug Therapeutics Program (DTP) AIDS Antiviral Screen (Zaharevitz, 2015), focuses on predicting the inhibition of HIV replication. The BACE dataset, included in MoleculeNet (Wu et al., 2018), measures the binding affinity of various inhibitors to $\beta$-secretase 1 (BACE-1). The overall dataset statistics of MoleculeNet are shown in Table  8.
@@
Table 8. Dataset statistics of MoleculeNet for molecule property prediction task.
@@
Table 9. Comparison of pre-train and inference times of FineMolTex, MolCa, and MoleculeSTM.
Device Pre-train Time Inference Time
MoleculeSTM 1 A100 40G 65h 47s
MolCA 2 A100 40G 33h 71s
FineMolTex 1 A100 40G 78h 89s
C.2. Baselines
@@ C.3. Operating Environment
C.4. Implementation Details
We use Pytorch to implement our model. For the pre-trained baselines, we directly utilize the checkpoints provided by the authors. For other baselines, we utilize the original codes from their authors and train the models in an end-to-end way. We utilize the BRICS (Degen et al., 2008) to fragment molecules into motifs. As BRICS primarily cleaves bonds according to a predefined set of chemical reactions, often resulting in several large fragments per molecule, we further utilize the post-processing procedure proposed by MGSSL (Zhang et al., 2021), which is designed to minimize the number of ring variants and facilitate the cleavage of side chains. Subsequently, we build a vocabulary of all motif tokens identified in the PubChemSTM dataset, which comprises a total of 30,080 unique motifs. We note that within our dataset, certain motifs are infrequently present, while others are prevalent but lack significant semantic value. This uneven distribution presents a challenge for the model, as it struggles to extract meaningful insights from such disparate occurrences. Thus we have constructed a masking set of 2,457 motifs that excludes motifs appearing fewer than 8 times or more than 80,005 times. Based on this masking set, we selectively mask motifs to enhance the model’s learning efficiency and focus on extracting valuable information from the most informative tokens. To map the embedding of the molecule and the text into the same dimension, we further utilize a projector MLP layer for the graph encoder. We utilize Adam as the optimizer for the GNN encoder, the language model, the two projector MLPs, the multi-modal framework, and the motif classifier with different learning rates, and test the learning rate ranging from {1e-3, 1e-4, 1e-5}. We set $\ell=2$, and test $\ell_{\text{trm}_{M}}$ raging from {1,2,3,4,5}, $\ell_{\text{trm}_{T}}$ ranging from {8,10,12}. For the coefficient of loss function, we test $\alpha$ and $\beta$ raging from {0.5,1,2}. We train our model with a total batch size of 16. We first train only the contrastive alignment task for 3 epochs, and then train both pretraining tasks together for 10 epochs. For fair comparisons, we randomly run 5 times and report the average results for all methods. The code, checkpoints, and optimal parameters can be found in the supplementary material and the anonymous repository https://anonymous.4open.science/status/FineMolTex-2266.
Optimal Parameters. The optimal parameters can be found in the config.json file in our code repository. Specifically, we set the dimension of the graph encoder to 300 and the dimensions of the transformer and cross-attention layer to 768. The hyperparameters are set as $\alpha=0.5$ and $\beta=1$. For $\ell=2$, we set $\ell_{\text{trm}M}$ to 2 for the first round and 3 for the second round, and $\ell_{\text{trm}_{T}}$ to 10 for the first round and 12 for the second round. The learning rates are set as follows: 1e-5 for the graph encoder, text encoder, transformer layers, and cross-attention layer; 5e-5 for the projector MLP of the graph encoder; and 1e-3 for the motif classifier.
@@ Appendix D Additional Resuts
D.1. Final Importance Scores
After pre-training, the final importance scores are shown in Table
Top 10 Bottom 10
Token Score Token Score
[NH4+] 0.2051 C=CC(C)C 0.0320
[NH3+] 0.1984 C 0.0315
Cl 0.1863 O 0.0282
[OH-] 0.1727 CCN(C)C 0.0278
C1=CC=CC=C1 0.1655 CCCCCNC 0.0263
CHO 0.1619 OCO 0.0236
Br 0.1561 COCOOC 0.0205
C1=CNC=C1 0.1522 CCCOCC 0.0174
F 0.1517 CCC(CC)OC 0.0157
[…]
Table 10. Top 10 and bottom 10 motif tokens by average importance score
Top 10 Bottom 10
Token Score Token Score
phosphatidyl 0.1623 cannot 0.002
amine 0.1602 participates 0.002
aromatic 0.1546 giving 0.002
nitrate 0.1492 mark 0.002
benzene 0.1405 eleven 0.002
sulphate 0.1378 year 0.002
ammonium 0.1362 errors 0.002
chloride 0.1351 particular 0.002
nitrate 0.1326 contributing 0.002
aldehyde 0.1284 affects 0.002
Table 11. Top 10 and bottom 10 word tokens by average importance score
D.2. Effectiveness of Importance-Based Selective Masking
We also trained “random mask only” for more epochs, extending it to 15 epochs. The results are shown in Table  LABEL:tab:effectiveness. We found that “random mask only” performance was still inferior to FineMolTex, which trained both pre-training tasks simultaneously for only 10 epochs. This highlights the efficiency of Selective Masking in accelerating learning and achieving better molecule-text alignment with fewer training steps.
Table 12. Ablation study (%±$\pm\sigma$) on molecule-ATC and DrugBank-Pharmacodynamics.
molecule-ATC DrugBank-Pharmacodynamics
Given Molecular Graph Given Text Given Molecular Graph Given Text
random mask only (15 epochs) 75.43±$\pm 0.15$ 75.22±$\pm 0.12$ 95.86±$\pm 0.34$ 95.80 ±$\pm 0.06$
FineMolTex 76.52±$\mathbf{\pm 0.10}$ 76.38±$\mathbf{\pm 0.18}$ 96.78±$\mathbf{\pm 0.05}$ 96.29±$\mathbf{\pm 0.12}$
D.3. Visualization of Motif and Word Tokens
Figure  8 displays the complete visualization with legends, illustrating that the embeddings of relevant motifs and words are closer in the embedding space.
Figure 8. Visualization of motif tokens and word tokens using t-SNE with legend.
D.4. Pre-training and inference times
We compare the pre-training and inference times of FineMolTex with MoleculeSTM and MolCA. For inference, we conduct experiments on the zero-shot graph-text retrieval task on DrugBank-Pharmacodynamics. Both MoleculeSTM and FineMolTex are tested on one NVIDIA A100 40 GB GPU, while MolCA is tested on two NVIDIA A100 40 GB GPUs. As shown in Table  LABEL:tab:time, the pre-training and inference times for FineMolTex are comparable to those of the baselines. Despite the longer pre-training time, we believe this trade-off is justified by the SOTA performance of FineMolTex in various downstream tasks.